OpenStudy (anonymous):

For this problem, the answer choices are "always", "sometimes", or "never". x + x = 2x For this problem, the answer choices are "always", "sometimes", or "never". x + x = 2x @Mathematics

OpenStudy (anonymous):

I do believe it would be always. ex) 1+1=2(1) --> 2=2 Because x+x=2x due to principles of algebra, you can say 2x=2x, which is always true.

OpenStudy (anonymous):

Thank you